OpenStudy (anonymous):

∫∫sin (xy+y) dA. where R= {(x,y)|0<=x<=pi, 0<=y<=pi/2 how to do the integration?

OpenStudy (amistre64):

is there a trig indentity for sin (a+b)?

OpenStudy (amistre64):

just an idea, dunno how useful it may be tho

OpenStudy (anonymous):

x and y are in cartesian coordinates? Just to be sure

OpenStudy (anonymous):

yes.

OpenStudy (amistre64):

sin (xy+y) = sin(xy)cos(y)+sin(y)cos(xy) this creates a constant in y so that you only have to deal with the x parts first

OpenStudy (anonymous):

but.....still confused any simple example

OpenStudy (amistre64):

\[\int_{0}^{pi/2}\int_{0}^{pi}sin (xy+y)~ dxdy\]\[\int_{0}^{pi/2}\int_{0}^{pi}sin(xy)cos(y)+sin(y)cos(xy)~ dxdy \] since y is a constant with respect to x .... \[\int_{0}^{pi/2}\int_{0}^{pi}sin(Cx)A+Bcos(Cx)~ dxdy \]

OpenStudy (amistre64):

\[\int_{0}^{pi/2}\int_{0}^{pi}sin(xy)cos(y)+sin(y)cos(xy)~ dxdy\] \[\int_{0}^{pi/2}\frac1ycos(\frac{pi}2 y)cos(y)+\frac1ysin(y)sin(\frac{pi}2y)-\frac1ycos(y)~dy\] might have kept track of all that

OpenStudy (anonymous):

is there any other method? method by substution ?

OpenStudy (amistre64):

there are always other methods .....

OpenStudy (anonymous):

ok. if possible able to let me knw..

OpenStudy (amistre64):

system froze; but i see a few errors; sin comes from -cos, and i used the wrong limits ....

OpenStudy (anonymous):

ok....

OpenStudy (amistre64):

so, focusing on the innards .... \[\int_{0}^{pi}sin(xy)cos(y)+sin(y)cos(xy)~ dx\] \[\left.-\frac1ycos(xy)cos(y)+\frac1ysin(y)sin(xy)\right|_{0}^{pi}\] \[-\frac1ycos(pi~y)cos(y)+\cancel{\frac1ysin(y)sin(pi~y)}^0+\frac1y\cancel{cos(0y)}^1 cos(y)-\cancel{\frac1ysin(y)sin(0y)}^0\] \[\frac1ycos(y)(1-cos(pi~y))\] hmmm

OpenStudy (amistre64):

maybe easier to rewrite it at this stage \[-\frac1y(cos(xy)cos(y)-sin(y)sin(xy))\] \[-\left.\frac1ycos(xy+y)\right|_{0}^{pi}\]

OpenStudy (amistre64):

\[−\frac1ycos((1+pi)~y)+\frac1ycos(y))\] \[\int_{0}^{pi/2}−\frac1ycos((1+pi)~y)+\frac1ycos(y))~dy\]

OpenStudy (anonymous):

ok.

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate%28integrate+sin%28xy%2By%29%2C+x%3D0..pi%29%2C+y%3D0..pi%2F2 ewww, that aint pretty no matter what

OpenStudy (amistre64):

i cant get a nice run on it ....

OpenStudy (anonymous):

its ok. thanks lot for the steps.

OpenStudy (amistre64):

good luck with it ;)

OpenStudy (amistre64):

also, make sure youve typed the question in correctly